Toxaway Company is a merchandiser that segments its business into two divisions-Commercial and Residential. The companys accRequired: 1. Do you agree with the interns decision to use an absorption format for her segmented income statement? 2. BasedReq 1 Req 2A Req 2B Req3 Req 4 Reg 5 Req 6 Req 7 Based on a review of the interns segmented income statement, how much of thReq 1 Req 2A Req 2B Req3 Req 4 Reg 5 Reg 6 Req 7 Based on a review of the interns segmented income statement, which of the fReal Req 1 Req 2A ReqzA Req 2B Req 28 | Req3 Req3 | Rega Req 4 Reas Req 5 Req 6 Rego Req 7 Rega Do you agree with the internReq 1 Req 2A Req 2B Req3 Req 4 Req 5 Req 6 Req 7 Redo the interns segmented income statement using the contribution format.Req 1 Req 2A Req 2B Req3 Req 4 Reg 5 Req 6 Req 7 Compute the companywide break-even point in dollar sales. (Round intermediatReg 1 Req2A Req 2A Req 2B Req2B Reg 3 Req3 Reg 4 Rega Req 5 Reqs | Rego Req 6 Req 7 Req7 Compute the break-even point in dollReq 1 Req 2B Req 2A Reg 3 Rego Reg? Req 4 Req 4 Regs Reg 5 Reg 6 Req 7 Assume the company decided to pay its sales representa
Answer #1
1.No, do not agree
2.a.
Allocation of common expenses is as follows
Commercial Residential
Selling and Admin Expenses 116,000 148,000
Less: Sales Commission 28,000 56,000
Less: Traceable fixed expenses 62,000 40,000
Common Expenses Allocated 26,000 52,000
b.Sales, is used since expenses allocated in the ratio of 1:2
3.No, Since common expenses are unavoidable
4.Contribution Margin format Income Statement
Total Commercial Residential
Sales 840,000 280,000 560,000
Less: Variable Expenses
Cost of Goods Sold 557,200 154,000 403,200
Selling and Admin Expenses 84,000 28,000 56,000
Contribution Margin 198,800 98,000 100,800
Less: Traceable Fixed Expenses 102,000 62,000 40,000
Segment Margin 96,800 36,000 60,800
Less: Common Fixed Expenses 78,000
Net Operating Income 18,800
5.Contribution Margin Ratio = Contribution Margin/Sales
=198800/840500
23.67%
Break even point = Fixed costs/CM Ratio
760563.3803
6. Commercial Division = 62,000/35% = $177,142.86
Residential = 40,000/18% = $222,222.22
7.Commercial = (62000+17000)/40% = $197500
Residential = (40,000+34,000)/23% = $321,739.13